A plan wave(500A•) falls normally on a long narrow slit of width 0.5mn. Calculate the angles of diffraction corresponding to the first three minima. Repeat the calculations crossponding width of 0.1mm. Interpred physically the change in the diffraction pattern.
1
Expert's answer
2020-07-29T09:48:57-0400

As per the question,

Wavelength of the plane wave "(\lambda)=500 A^\circ =500\times 10^{-10}m"

Width of the slit "=0.5 mm = 5\times 10^{-4}m"

Angle of diffraction = ?

"\theta =\frac{\lambda}{w}"

"\theta = \frac{500 \times 10^{-10}m}{5\times 10^{-4}m} =1\times 10^{-4} radian"

similarly,

"\theta' = \frac{500 \times 10^{-10}m}{1\times 10^{-4}m} =5\times 10^{-4} radian"